# A geometric approach to exponentially small splitting: The generic zero-Hopf bifurcation of co-dimension two

### 12 Mar 2026 v1

@@ 1. Introduction
The proof of the expression (1.7) in [1] follows a well-known approach (first initiated by Lazutkin for maps, see [32]): First, the authors define $(x_{2},y_{2},z_{2})$ by
@@
which brings (1.5) into the following slow-fast form
@@
with $t_{2}=\mathcal{O}(1)$, a so-called inner problem is defined for $\epsilon\to 0$, and it is shown that the unstable and stable manifolds lie close to special invariant manifold solutions of this inner problem. In this way, given that the distance between the invariant manifolds is known near the poles (with $t_{2}$ fixed, see (1.11)), the desired separation of the unstable and stable manifolds at $t=0$ (corresponding to $x=0$ by (1.10)) is obtained through the solution of a certain linear boundary value problem (obviously up to many technical details; see also [21, pp. 673–674] for a related summary of the approach).
The constant $C$ in (1.7) (known as a Stokes constant) is essentially a measure of the separation of the invariant manifold solutions of the inner problem.
The approach in [1], which is more functional-analytic in nature than dynamical-systems-oriented, has been successful in many important problems with exponentially small splitting, see e.g. [2, 4, 18, 21] and references therein. However, the point of departure (also within the context of maps [19, 32]) is always a parametrization of the unperturbed solution and its singularities in the complex plane, which is not only a strong assumption in a general context, but also an unusual one from the perspective of dynamical systems theory.
@@ 1.1. Main results
In this paper, we provide an alternative method for the asymptotic splitting of the one-dimensional invariant manifolds. In particular, our approach does not use complex time and the explicit solution of the $\epsilon=0$ limit, but instead we work exclusively in the complexified phase space $(x,y,z)\in\mathbb{C}^{3}$, parametrizing our invariant manifolds as graphs over $x$. Our approach is therefore more in tune with dynamical systems theory. Moreover, it is our anticipation that the method will be useful in problems where explicit time parametrizations and their singularities are unknown.
@@
The unperturbed invariant manifolds $(y,z)=\psi_{0}^{\pm}(x)$, $x\in S^{\pm}$, are obtained by Borel-Laplace. In particular, the manifolds are analytic if and only if the Borel transform $\mathcal{B}(\psi_{0})$ defines an entire function with at most exponential growth, see [6, 13]. However, the linear part of the equation for $\mathcal{B}(\psi_{0})$ has poles at $\pm i$ and this is the general case for the Borel transform of $\psi_{0}^{\pm}$ too, see also [11, 28, 34, 40].
@@
would be a more general statement for the splitting, with $\widetilde{\phi}_{0}(0)\neq 0$ if the unperturbed manifolds are nonanalytic at the origin (this will be a simple consequence of (5.26) below). More generally, we have
@@
where $\beta\in\mathbb{N}$ is the smallest number for which the higher order “corrections” $\psi_{1,\beta}^{\pm}$, to the unperturbed manifold, are nonanalytic at the origin, see Lemma 4.3. We therefore speculate whether the following holds true: The formal series expansion of the connection converges or there is some $\beta\in\mathbb{N}$ and some $C_{\beta}\neq 0$ such that
@@
We leave this open to future work.
@@ 1.2. A summary of our approach
to extend the unstable and stable manifolds, which are naturally parameterized in compact subsets of the $(x_{2},y_{2},z_{2})$-space as graphs $(y_{2},z_{2})=m_{2}^{\pm}(x_{2},\epsilon)$ over $x_{2}\in\mathcal{X}_{2}^{\pm}$, see (1.9), to order $x=\mathcal{O}(1)\in\mathbb{C}$, $|x|>0$ sufficiently small. For the extension, we apply the flow in appropriate normal form coordinates, obtained by following the approach in [26], using two separate notions of formal invariant manifolds: (i) formal series with respect to $r_{1}$ with $\epsilon_{1}$-dependent coefficients, and (ii) formal series with respect to $\epsilon_{1}$ with $r_{1}$-dependent coefficients (with $r_{1}\in S^{\pm}\subset\mathbb{C}$). Our result shows that the unperturbed manifolds give the leading order approximation of the unstable and stable manifolds $(y,z)=m^{\pm}(x,\epsilon)$, for $x\in S^{\pm}$, $|x|>0$ small enough, respectively, as $\epsilon\to 0$. Therefore, whenever the unperturbed manifolds are distinct on the common domain $S^{+}\cap S^{-}$, we obtain an $\mathcal{O}(1)$-separation of the unstable and stable manifolds for $x\in S^{+}\cap S^{-}$, $|x|\geq\xi>0$ small enough, for $0<\epsilon\ll 1$. To determine the separation of the invariant manifolds at $\{x=0\}$, we then derive a linear equation for the difference $(\Delta y,\Delta z)=(m^{+}-m^{-})(x,\epsilon)$ (as in [1]) for $x$ on the imaginary axis $x\in i\mathbb{R}$, $|x|\geq 0$ small enough. As a novel aspect, we identify this set of equations as a slow-fast system with a normally hyperbolic critical manifold (of saddle-type) for $\epsilon\to 0$. Therefore, in contrast to the functional-analytic approach of [1], we use Geometric Singular Perturbation Theory (GSPT) [16, 25] to study this linear system and derive a Fenichel-type normal form [25, Section 5], bringing the linear system to diagonal form. In this way, the separation can easily be integrated from $x=\pm i\delta$, $\delta>0$ fixed small enough, to $x=0$. Here we also use that the equations are real-analytic: $(x,y,z)(t)$ is a solution of (1.5), $\epsilon>0$, if only if $(\overline{x},\overline{y},\overline{z})(t)$ is a solution, where the bar denotes the complex conjugate. We illustrate our approach in Fig. 2.
@@ 2. The $\epsilon=0$-system: The unperturbed invariant manifolds
This system is a generalized saddle-node with Poincaré rank $1$ studied by many authors, see e.g. [6]. In particular, the linearization of the right hand side for $r_{1}=0$ at $\zeta_{1}=(0,0)$ is given by $\Omega$, having nonzero eigenvalues $\pm i$. It then follows from [6, Theorem 3] (as a corollary of the invariance of $y=0$ for the normal form [6, Eq. (8)], see also [26, Lemma 4.8] or [13, Corollary 4.2] where the statement is made more explicit) that there exists two solutions
@@ 3. The perturbed system: The unstable and stable manifolds
By (3.2), we have that for any $\varrho>0$, there is an $\epsilon_{0}>0$ such that $F_{2}$, $G_{2}$ and $H_{2}$ are real-analytic on $(x_{2},y_{2},z_{2},\epsilon)\in B_{\varrho}^{3}\times(-\epsilon_{0},\epsilon_{0})$. Here $B_{\varrho}^{3}\subset\mathbb{C}^{3}$ is the ball of radius $\varrho>0$ centered at the origin. But then [12, Proposition 2.1] gives the existence and uniqueness of the formal series (3.8) with $m_{2,\alpha}:B_{\varrho}\rightarrow\mathbb{C}^{2}$ real-analytic for all $\alpha\in\mathbb{N}$ and any $\varrho>0$. The result therefore follows. ∎
@@
for $t\geq 0$, see e.g. [35, Theorem 5.3]. Here $\widetilde{\mathcal{M}}_{2}^{N}(t,\widetilde{x}_{20},\epsilon)$ is the fundamental matrix associated with the linear problem
@@
see e.g. [10, Proposition 1, p. 34], for $c_{1}>0,c_{2}>0$ both small enough, for any $s\geq t$, $\widetilde{x}_{2}\in\widetilde{\mathcal{X}}_{2}^{-}$, and all $0<\epsilon\ll 1$. Here $\|\cdot\|$ denotes the operator norm. To obtain (3.18), we have also used that $\|\exp({\epsilon^{-1}\Omega t})\|\equiv 1$, recall (2.6).
The existence of a unique bounded solution $t\mapsto\widetilde{\zeta}_{2}(t,\widetilde{x}_{20},\epsilon),t\geq 0$, of the fixed-point equation (3.16), then proceeds completely analogously to [35, Theorem 5.3] for any $0<\epsilon\ll 1$. The desired graph is given by $\widetilde{\zeta}_{2}=\widetilde{\zeta}_{2}(0,\widetilde{x}_{2},\epsilon)=\mathcal{O}(\epsilon^{N})$, uniformly over $\widetilde{x}_{2}\in\widetilde{\mathcal{X}}_{2}^{-}$, for all $0\leq\epsilon\ll 1$. Next, for the smoothness with respect to $\epsilon$, we first note that the stable manifold is unique, being independent of $N$, and smooth (even analytic) for $\epsilon\in(0,\epsilon_{0})$ by standard theory. The boundedness of $\widetilde{\zeta}_{2}$ for any $N$ and $0<\epsilon\ll 1$ then implies (by (3.14)) that the invariant manifold has an asymptotic expansion with respect to $\epsilon\to 0$, uniformly with respect to $x_{2}\in\mathcal{X}_{2}^{-}$. Finally, for any $k\in\mathbb{N}$ there is an $N\in\mathbb{N}$ ($N\geq 2k+1$ will do) such that the right hand side of (3.16) is $C^{k}$-smooth with respect to $\epsilon\in[0,\epsilon_{0})$; notice that we lose two powers of $\epsilon$ upon differentiation due to (3.17). It follows that the invariant manifold is $C^{\infty}$-smooth with respect to $\epsilon\in[0,\epsilon_{0})$. ∎
@@ 4.1. Analysis in the $\breve{x}=1$-chart
in the formal sense. In particular, we will (as in [26, Section 4.3]) use two different notions of formal series: Formal series in $r_{1}$ with $\epsilon_{1}$-depending coefficients:
@@
for some locally defined real-analytic function $R_{1}$. This follows from a simple calculation. The result can then be proven as the proof of [12, Proposition 2.1.] (by the virtue of the $r_{1}$-factor on the right hand side). ∎
@@
Moreover, we expand
$(1-\epsilon_{1}^{2})^{-1}$ $=:\sum_{\alpha=0}^{\infty}A_{1,\alpha}\epsilon_{1}^{\alpha},\quad A_{1,\alpha}=\begin{cases}1&\mbox{$\alpha$ even},\\ 0&\mbox{$\alpha$ odd},\end{cases}$
$R_{1}(r_{1},\zeta_{1},\epsilon_{1})$ $=:\sum_{\alpha,\beta}R_{1,\alpha,\beta}(r_{1})\zeta_{1}^{\alpha}\epsilon_{1}^{\beta},$
as convergent series, and with $\zeta_{1}$ as the formal series (4.13) (for simplicity we drop the $\pm$-superscript, which refers to the direction $S^{\pm}$), we define $(\zeta_{1}^{\alpha})_{\gamma}$, $\alpha\in\mathbb{N}$, by
$\zeta_{1}^{\alpha}=:\sum_{\gamma=0}^{\infty}(\zeta_{1}^{\alpha})_{\gamma}\epsilon_{1}^{\gamma}.$
We suppress the $r_{1}$-dependency of $(\zeta_{1}^{\alpha})_{\gamma}$ in the following. By Cauchy’s product rule and the multinomial theorem, $(\zeta_{1}^{\alpha})_{\gamma}=(\zeta_{1}^{\alpha})_{\gamma}(\psi_{1,0},\ldots,\psi_{1,\gamma})$ is a homogeneous polynomial of degree $\alpha\in\mathbb{N}$ in the variables $\psi_{1,0},\ldots,\psi_{1,\gamma}$ and given by
$(\zeta_{1}^{\alpha})_{\gamma}(\psi_{1,0},\ldots,\psi_{1,\gamma})=\sum_{\upsilon}\begin{pmatrix}\alpha\\ \upsilon_{0},\,\upsilon_{1},\ldots,\upsilon_{\gamma}\end{pmatrix}\psi_{1,0}^{\upsilon_{0}}\psi_{1,1}^{\upsilon_{1}}\cdots\psi_{1,\gamma-1}^{\upsilon_{\gamma-1}}\psi_{1,\gamma}^{\upsilon_{\gamma}}.$ (4.15)
Here
$\begin{pmatrix}\alpha\\ \upsilon_{0},\upsilon_{1},\ldots,\upsilon_{\gamma-1},\upsilon_{\gamma}\end{pmatrix}=\frac{\alpha!}{\upsilon_{0}!\upsilon_{1}!\cdots\upsilon_{\gamma-1}!\upsilon_{\gamma}!},$
denotes the multinomial coefficient and the sum in (4.15) runs over all $\upsilon=(\upsilon_{0},\ldots,\upsilon_{\gamma})\in\mathbb{N}_{0}^{\gamma+1}$ such that $\sum_{\iota=0}^{\gamma}\upsilon_{\iota}=\alpha$ and $\sum_{\iota=0}^{\gamma}\iota\upsilon_{\iota}=\gamma$. Notice that
$(\zeta_{1}^{\alpha})_{\gamma}(\psi_{1,0},\ldots,\psi_{1,\gamma})=\alpha\psi_{1,0}^{\alpha-1}\psi_{1,\gamma}+M_{1,\alpha,\gamma-1}(\psi_{1,0},\ldots,\psi_{1,\gamma-1})$
where $M_{1,\alpha,\gamma-1}$ is given by
$M_{1,\alpha,\gamma-1}(\psi_{1,0},\ldots,\psi_{1,\gamma-1})=\sum_{\upsilon}\begin{pmatrix}\alpha\\ \upsilon_{0},\,\upsilon_{1},\ldots,\upsilon_{\gamma-1}\end{pmatrix}\psi_{1,0}^{\upsilon_{0}}\psi_{1,1}^{\upsilon_{1}}\cdots\psi_{1,\gamma-1}^{\upsilon_{\gamma-1}},$
for all $\alpha\in\mathbb{N}_{0},\gamma\in\mathbb{N}$. The sum now runs over all $\upsilon=(\upsilon_{0},\ldots,\upsilon_{\gamma-1})\in\mathbb{N}_{0}^{\gamma}$ such that $\sum_{\iota=0}^{\gamma-1}\upsilon_{\iota}=\alpha$ and $\sum_{\iota=0}^{\gamma-1}\iota\upsilon_{\iota}=\gamma$. Consequently, by Cauchy’s product rule
$R_{1}(r_{1},\zeta_{1},\epsilon_{1})$ $=R_{1}(r_{1},\psi_{1,0},0)+\sum_{\gamma=1}^{\infty}\Bigg[\sum_{\alpha=0}^{\infty}\Bigg(R_{1,\alpha,0}\alpha\psi_{1,0}^{\alpha-1}\psi_{1,\gamma}$
$+R_{1,\alpha,0}M_{1,\alpha,\gamma-1}+\sum_{\beta=0}^{\gamma-1}R_{1,\alpha,\gamma-\beta}(\zeta_{1}^{\alpha})_{\beta}\Bigg)\Bigg]\epsilon_{1}^{\gamma}$
$=R_{1}(r_{1},\psi_{1,0},0)+\sum_{\gamma=1}^{\infty}\Bigg[\frac{\partial R_{1}}{\partial\zeta_{1}}(r_{1},\psi_{1,0},0)\psi_{1,\gamma}$
$+\sum_{\alpha=0}^{\infty}\Bigg(R_{1,\alpha,0}M_{1,\alpha,\gamma-1}+\sum_{\beta=0}^{\gamma-1}R_{1,\alpha,\gamma-\beta}(\zeta_{1}^{\alpha})_{\beta}\Bigg)\Bigg]\epsilon_{1}^{\gamma}$
In this way, we obtain the following equations for $\psi_{1,\gamma}$, $\gamma\in\mathbb{N}_{0}$:
@@
$r_{1}^{2}\psi_{1,\gamma}^{\prime}$ $=\Omega\psi_{1,\gamma}+r_{1}\left(\frac{\partial R_{1}}{\partial\zeta_{1}}(r_{1},\psi_{1,0},0)+\gamma\operatorname{Id}\right)\psi_{1,\gamma}$ (4.16)
$+Q_{1,\gamma-1}(r_{1},\psi_{1,0},\ldots,\psi_{1,\gamma-1})\quad\forall\,\gamma\in\mathbb{N},$
@@
$Q_{1,\gamma-1}(r_{1},\psi_{1,0},\ldots,\psi_{1,\gamma-1})$ $:=\sum_{\alpha=0}^{\gamma-1}\Omega A_{1,\gamma-\alpha}\psi_{1,\alpha}$
$+r_{1}\sum_{\alpha=0}^{\infty}\Bigg(R_{1,\alpha,0}M_{1,\alpha,\gamma-1}+\sum_{\beta=0}^{\gamma-1}R_{1,\alpha,\gamma-\beta}(\zeta_{1}^{\alpha})_{\beta}\Bigg),$
is a locally defined real-analytic function. It follows that $\psi_{1,\gamma}\in S^{\pm}(\delta_{\gamma},\chi)$ for $\delta_{\gamma}>0$ sufficiently small for all $\gamma\in\mathbb{N}_{0}$ from [6] (as in the proof of Lemma 2.1). In particular, $\psi_{1,\alpha}^{\pm}$ are the $1$-sums of a Gevrey-$1$ series in the directions defined by $S^{\pm}$, respectively. To obtain that $\delta_{\gamma}>0$ is uniformly bounded from below, we use that (4.16) is a linear first order equation for $\psi_{1,\gamma}$, $\gamma\in\mathbb{N}$, which is regular for $r_{1}\neq 0$. Therefore, by Lemma 2.1 ($\gamma=0$ case), the existence and uniqueness of solutions of linear equations and induction on $\gamma$, we conclude that $\psi_{1,\gamma}$ can be extended to $S^{\pm}(\delta,\chi)$, $\delta=\delta_{0}$, for all $\gamma\in\mathbb{N}_{0}$. ∎
@@
bring (4.8) into the following two sets of equations
@@
We follow [26, Lemma 4.5] and take

### 17 Apr 2026 v2

@@ 1. Introduction
The proof of the expression (1.7) in [1] follows a well-known approach (initiated by Lazutkin for maps, see [35]): First, the authors scale the variables. In the present paper, we will use the following scaling
@@
rather than the scaling used in [1]
$\begin{cases}x=\epsilon x_{2},\\ y=\epsilon y_{2},\\ z=\epsilon z_{2},\end{cases}$
but this difference is not essential. The transformation defined by (1.8) brings (1.5) into the following slow-fast form
@@
with $t_{2}=\mathcal{O}(1)$, a so-called inner problem is defined for $\epsilon\to 0$, and it is shown that the unstable and stable manifolds lie close to special invariant manifold solutions of this inner problem. In this way, given that the distance between the invariant manifolds is known near the poles (with $t_{2}$ fixed, see (1.11)), the desired separation of the unstable and stable manifolds at $t=0$ (corresponding to $x=0$ by (1.10)) is obtained through the solution of a certain linear boundary value problem (obviously up to many technical details; see also [24, pp. 673–674] for a related summary of the approach).
The constant $C_{0}$ in (1.7) (known as a Stokes constant) is essentially a measure of the separation of the invariant manifold solutions of the inner problem.
The approach in [1], which is more functional-analytic in nature than dynamical-systems-oriented, has been successful in many important problems with exponentially small splitting, see e.g. [2, 3, 4, 6, 21, 24] and references therein. However, the point of departure (also within the context of maps [22, 35]) is always a parametrization of the unperturbed solution and its singularities in the complex plane, which is not only a strong assumption in a general context, but also an unusual one from the perspective of dynamical systems theory.
@@ 1.1. Main results
In this paper, we provide an alternative method for the asymptotic splitting of the one-dimensional invariant manifolds. In particular, our approach does not use complex time and the explicit solution of the $\epsilon=0$ limit, but instead we work exclusively in the complexified phase space $(x,y,z)\in\mathbb{C}^{3}$, parametrizing our invariant manifolds as graphs over $x$. Our approach is therefore more in tune with dynamical systems theory. It is our anticipation that the method will be useful in problems where explicit time parametrizations and their singularities are unknown.
@@
The unperturbed invariant manifolds $(y,z)=\psi_{0}^{\pm}(x)$, $x\in S^{\pm}$, are obtained by Borel-Laplace. In particular, the manifolds are analytic if and only if the Borel transform $\mathcal{B}(\psi_{0})(w)$ defines an entire function with at most exponential growth for $w\rightarrow\infty$, see [8, 16]. However, the linear part of the equation for $\mathcal{B}(\psi_{0})(w)$ has poles at $w=\pm i$ and this is also the general case for the actual Borel transform of $\psi_{0}^{\pm}$ (for the full nonlinear equation, see also [14, 31, 37, 43]).
@@
would be a more general statement for the splitting, with $\widetilde{\phi}_{0}(0)\neq 0$ if the unperturbed manifolds are nonanalytic at the origin. More generally, we have
@@
provided that $\beta\in\mathbb{N}$ is the smallest number for which the higher order “corrections” $\psi_{1,\beta}^{\pm}$, to the unperturbed manifold, are nonanalytic at the origin, see Lemma 4.3. These statements are simple consequences of (5.26) below. We therefore speculate whether the following holds true: Either
  1. (1)
    the formal series expansion of the connection converges and $d(\epsilon)=0$ for all $0<\epsilon\ll 1$,
or
  1. (2)
    there is some $\beta\in\mathbb{N}_{0}$ and some $C_{\beta}\neq 0$ such that
@@
This conjecture relates to the existence of “ultra-exponentially small splitting of separatrices” as discussed in [20]. We leave this open to future work.
@@ 1.2. A summary of our approach
to extend the unstable and stable manifolds, which are naturally parameterized in compact subsets of the $(x_{2},y_{2},z_{2})$-space as graphs $(y_{2},z_{2})=m_{2}^{\pm}(x_{2},\epsilon)$ over $x_{2}\in\mathcal{X}_{2}^{\pm}$, see (1.9), to order $x=\mathcal{O}(1)\in\mathbb{C}$, $|x|>0$ sufficiently small. For the extension, we apply the flow in appropriate normal form coordinates, obtained by following the approach in [29], using two separate notions of formal invariant manifolds: (i) formal series with respect to $r_{1}$ with $\epsilon_{1}$-dependent coefficients, and (ii) formal series with respect to $\epsilon_{1}$ with $r_{1}$-dependent coefficients (with $r_{1}\in S^{\pm}\subset\mathbb{C}$). Our result shows that the unperturbed manifolds give the leading order approximation of the unstable and stable manifolds $(y,z)=m^{\pm}(x,\epsilon)$, for $x\in S^{\pm}$, $|x|>0$ small enough, respectively, as $\epsilon\to 0$. Therefore, whenever the unperturbed manifolds are distinct on the common domain $S^{+}\cap S^{-}$, we obtain an $\mathcal{O}(1)$-separation of the unstable and stable manifolds for $x\in S^{+}\cap S^{-}$, $|x|\geq\xi>0$ small enough, for $0<\epsilon\ll 1$. To determine the separation of the invariant manifolds at $\{x=0\}$, we then derive a linear equation for the difference $(\Delta y,\Delta z)=(m^{+}-m^{-})(x,\epsilon)$ (as in [1]) for $x$ on the imaginary axis $x\in i\mathbb{R}$, $|x|\geq 0$ small enough. As a novel aspect, we identify this set of equations as a slow-fast system with a normally hyperbolic critical manifold (of saddle-type) for $\epsilon\to 0$. Therefore, in contrast to the functional-analytic approach of [1], we use Geometric Singular Perturbation Theory (GSPT) [19, 28] to study this linear system and derive a Fenichel-type normal form [28, Section 5], bringing the linear system to diagonal form. In this way, the separation can easily be integrated from $x=\pm i\delta$, $\delta>0$ fixed small enough, to $x=0$. Here we also use that the equations are real-analytic: $(x,y,z)(t)$ is a solution of (1.5), $\epsilon>0$, if only if $(\overline{x},\overline{y},\overline{z})(t)$ is a solution, where the bar denotes the complex conjugate. We illustrate our approach in Fig. 2.
@@ 2. The $\epsilon=0$-system: The unperturbed invariant manifolds
This system is a generalized saddle-node with Poincaré rank $1$ studied by many authors, see e.g. [8]. In particular, the linearization of the right hand side for $r_{1}=0$ at $\zeta_{1}=(0,0)$ is given by $\Omega$, having nonzero eigenvalues $\pm i$. It then follows from [8, Theorem 3] (as a corollary of the invariance of $y=0$ for the normal form [8, Eq. (8)], see also [29, Lemma 4.8] or [16, Corollary 4.2] where the statement is made more explicit) that there exists two solutions
@@ 3. The perturbed system: The unstable and stable manifolds
By (3.2), we have that for any $\varrho>0$, there is an $\epsilon_{0}>0$ such that $F_{2}$, $G_{2}$ and $H_{2}$ are real-analytic on $(x_{2},y_{2},z_{2},\epsilon)\in B_{\varrho}^{3}\times(-\epsilon_{0},\epsilon_{0})$. Here $B_{\varrho}^{3}\subset\mathbb{C}^{3}$ is the ball of radius $\varrho>0$ centered at the origin. But then [15, Proposition 2.1] gives the existence and uniqueness of the formal series (3.8) with $m_{2,\alpha}:B_{\varrho}\rightarrow\mathbb{C}^{2}$ real-analytic for all $\alpha\in\mathbb{N}$ and any $\varrho>0$. The result therefore follows. ∎
@@
for $t\geq 0$, see e.g. [38, Theorem 5.3]. Here $\widetilde{\mathcal{M}}_{2}^{N}(t,\widetilde{x}_{20},\epsilon)$ is the fundamental matrix associated with the linear problem
@@
see e.g. [13, Proposition 1, p. 34], for $c_{1}>0,c_{2}>0$ both small enough, for any $s\geq t$, $\widetilde{x}_{2}\in\widetilde{\mathcal{X}}_{2}^{-}$, and all $0<\epsilon\ll 1$. Here $\|\cdot\|$ denotes the operator norm. To obtain (3.18), we have also used that $\|\exp({\epsilon^{-1}\Omega t})\|\equiv 1$, recall (2.6).
The existence of a unique bounded solution $t\mapsto\widetilde{\zeta}_{2}(t,\widetilde{x}_{20},\epsilon),t\geq 0$, of the fixed-point equation (3.16), then proceeds completely analogously to [38, Theorem 5.3] for any $0<\epsilon\ll 1$. The desired graph is given by $\widetilde{\zeta}_{2}=\widetilde{\zeta}_{2}(0,\widetilde{x}_{2},\epsilon)=\mathcal{O}(\epsilon^{N})$, uniformly over $\widetilde{x}_{2}\in\widetilde{\mathcal{X}}_{2}^{-}$, for all $0\leq\epsilon\ll 1$. Next, for the smoothness with respect to $\epsilon$, we first note that the stable manifold is unique, being independent of $N$, and smooth (even analytic) for $\epsilon\in(0,\epsilon_{0})$ by standard theory. The boundedness of $\widetilde{\zeta}_{2}$ for any $N$ and $0<\epsilon\ll 1$ then implies (by (3.14)) that the invariant manifold has an asymptotic expansion with respect to $\epsilon\to 0$, uniformly with respect to $x_{2}\in\mathcal{X}_{2}^{-}$. Finally, for any $k\in\mathbb{N}$ there is an $N\in\mathbb{N}$ ($N\geq 2k+1$ will do) such that the right hand side of (3.16) is $C^{k}$-smooth with respect to $\epsilon\in[0,\epsilon_{0})$; notice that we lose two powers of $\epsilon$ upon differentiation due to (3.17). It follows that the invariant manifold is $C^{\infty}$-smooth with respect to $\epsilon\in[0,\epsilon_{0})$. ∎
@@ 4.1. Analysis in the $\breve{x}=1$-chart
in the formal sense. In particular, we will (as in [29, Section 4.3]) use two different notions of formal series: Formal series in $r_{1}$ with $\epsilon_{1}$-depending coefficients:
@@
for some locally defined real-analytic function $R_{1}$. This follows from a simple calculation. The result can then be proven as the proof of [15, Proposition 2.1.] (by the virtue of the $r_{1}$-factor on the right hand side). ∎
@@
With $\zeta_{1}$ as the formal series (4.13) (for simplicity we drop the $\pm$-superscript, which refers to the direction $S^{\pm}$), we use the Faa di Bruno formula [12, Theorem 2.1] (with $F(\zeta_{1},\epsilon_{1})=R_{1}(r_{1},\zeta_{1},\epsilon_{1})$, $G(\epsilon_{1})=(\zeta_{1}(\epsilon_{1}),\epsilon_{1})$) to write
$R_{1}(r_{1},\zeta_{1},\epsilon_{1})$ $=R_{1}(r_{1},\psi_{1,0},0)+\sum_{\alpha=1}^{\infty}R_{1,\alpha}(r_{1},\psi_{1,0},\ldots,\psi_{1,\alpha})\epsilon_{1}^{\alpha},$ (4.15)
where
$R_{1,\alpha}(r_{1},\psi_{1,0},\ldots,\psi_{1,\alpha})=$ $\sum_{1\leq|\lambda|\leq\alpha}\frac{1}{\lambda!}\frac{\partial^{|\lambda|}}{\partial\zeta_{1}^{(\lambda_{1},\lambda_{2})}\partial\epsilon_{1}^{\lambda_{3}}}R_{1}(r_{1},\psi_{1,0},0)\times$ (4.16)
$\sum_{\gamma=1}^{\alpha}\sum_{(\sigma,\ell)\in p_{\gamma,\alpha,\lambda}}\prod_{\beta=1}^{\gamma}\frac{1}{\sigma_{\beta}!(\ell_{\beta}!)^{|\sigma_{\beta}|}}(\ell_{\beta}!\psi_{1,\ell_{\beta}})^{(\sigma_{1\beta},\sigma_{2\beta})}(\delta_{1\beta})^{\sigma_{3\beta}}.$
Here $\delta_{\alpha\beta}$ denotes Kronecker’s delta. The details of the finite index sets
$p_{\gamma,\alpha,\lambda}\subset\left\{(\sigma,\ell)\in(\mathbb{N}^{3})^{\gamma}\times\mathbb{N}^{\gamma}\,:\,\sum_{\beta=0}^{\gamma}\sigma_{\beta}=\lambda\,\,\text{and}\,\,\sum_{\beta=0}^{\gamma}|\sigma_{\beta}|\ell_{\beta}=\alpha\right\}.$ (4.17)
are not important, and we therefore refer to [12, Eq. (2.2)]. We will only use that
$R_{1,\alpha}(r_{1},\psi_{1,0},\ldots,\psi_{1,\alpha}):=Q_{1,\alpha}(r_{1},\psi_{1,0},\ldots,\psi_{1,\alpha-1})+\frac{\partial}{\partial\zeta_{1}}R_{1}(r_{1},\psi_{1,0},0)\psi_{1,\alpha},$
with the last term obtained from the sum in (4.16) with $\lambda=(1,0,0)$ and $\lambda=(0,1,0)$ (where $p_{\gamma,\alpha,\lambda}$ are singletons), so that $Q_{1,\alpha}$ (cf. (4.16) and (4.17)) is a multivariate polynomial of degree $\alpha$ with respect to the variables
$(\psi_{1,1},\ldots,\psi_{1,\alpha-1}),$
for fixed $r_{1},\psi_{1,0}$, for all $\alpha\in\mathbb{N}$. In this way, by inserting (4.13) into (4.14) and collecting terms, we obtain the following equations for $\psi_{1,\alpha}$, $\alpha\in\mathbb{N}_{0}$:
@@
$r_{1}^{2}\psi_{1,\alpha}^{\prime}$ $=\Omega\psi_{1,\alpha}+r_{1}\left(\frac{\partial R_{1}}{\partial\zeta_{1}}(r_{1},\psi_{1,0},0)+\alpha\operatorname{Id}\right)\psi_{1,\alpha}$ (4.18)
$+\sum_{\gamma=0}^{\alpha-1}\Omega A_{1,\alpha-\gamma}\psi_{1,\gamma}+r_{1}Q_{1,\alpha-1}(r_{1},\psi_{1,0},\ldots,\psi_{1,\alpha-1})\quad\forall\,\alpha\in\mathbb{N},$
@@
$A_{1,\alpha}=\begin{cases}1&\mbox{$\alpha$ even},\\ 0&\mbox{$\alpha$ odd}.\end{cases}$
It follows that $\psi_{1,\alpha}\in S^{\pm}(\delta_{\alpha},\chi)$ for $\delta_{\alpha}>0$ sufficiently small for all $\alpha\in\mathbb{N}_{0}$ from [8] (as in the proof of Lemma 2.1). In particular, $\psi_{1,\alpha}^{\pm}$ are the $1$-sums of a Gevrey-$1$ series in the directions defined by $S^{\pm}$, respectively. To obtain that $\delta_{\alpha}>0$ is uniformly bounded from below, we use that (4.18) is a linear first order equation for $\psi_{1,\alpha}$, $\alpha\in\mathbb{N}$, which is regular for $r_{1}\neq 0$. Therefore, by Lemma 2.1 ($\alpha=0$ case), the existence and uniqueness of solutions of linear equations and induction on $\alpha$, we conclude that $\psi_{1,\alpha}$ can be extended to $S^{\pm}(\delta,\chi)$, $\delta=\delta_{0}$, for all $\alpha\in\mathbb{N}_{0}$. ∎
@@
bring (4.8) into the following two sets of equations by pull-back
@@
We follow [29, Lemma 4.5] and take